OpenStudy (anonymous):

y=e^x+sin(x) How can you show that this equation has real roots? (Maybe with rolle?)

OpenStudy (kc_kennylau):

Roots occur when y=0, the equation transforms to e^x+sin(x)=0

OpenStudy (anonymous):

Yes, but how can I proove that there are x's which solve this solution? :)

OpenStudy (anonymous):

*this equation

OpenStudy (kc_kennylau):

Find y when x=-pi/2 and when x=0 :)

OpenStudy (kc_kennylau):

y changes sign between x=-pi/2 and x=0, therefore a root must exist in (-pi/2,0) because this function is continuous

OpenStudy (kc_kennylau):

And this function is continuous because e^x and sin(x) are both continuous

OpenStudy (kc_kennylau):

Because if y changes sign in x=[-pi/2,0], y must pass through 0.

OpenStudy (anonymous):

that gives me 1 root (with the mean value theorem), but there is an infinite number of roots, right? (because of sin)

OpenStudy (kc_kennylau):

Yes, but this is out of the question's scope :)